If Josh spent half of his money, then he earned $6.00 and then lost 1/6 of his money and now he has $15.00. How much did he start with? Please show work.
Answer by stanbon(75887)

If Josh spent half of his money, then he earned $6.00 and then lost 1/6 of his money and now he has $15.00. How much did he start with?
---------------------
Let the amount he started with be "x".
Equation:
(1/2)x + 6 - (1/6)[(1/2)x+6] = 15
---
(5/6)[(1/2)x+6] = 15
(1/2)x + 6 = (6/5)*15
(1/2)x + 6 = 18
(1/2)x = 12
x = $24
=============
Cheers,
Stan H.